an alloy of gold and copper weighing 100 gm contains 97 gm of gold. how much gold should be added to the alloy to increase the percentage og gold to 98?
Answer by richwmiller(17219)

(97+x)/(100+x)=.98
97+x=.98*(100+x)
97+x=98+.98x
.02x=1
2x=100
x=50
check
97+50/150=147/150=,98
ok